OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (3, 6) and (–2, 1)?

OpenStudy (anonymous):

(y-y./x-x.) u got try

OpenStudy (anonymous):

\[\large \frac{y - y_1}{y_2 - y_1} = \frac{x - x_1}{x_2 - x_1}\] Put the values and then find the equation..

OpenStudy (anonymous):

my slope is one I know that and I use this formula y=mx=b 6=1(3)+b 6=3+b 6-3=b but there I got stuck

OpenStudy (anonymous):

draw like a chart and follow the slope pattern and look for your y intercept.. i don't know if that will work though

OpenStudy (anonymous):

so i got the y intercept as 3

OpenStudy (anonymous):

ok thanks ! :)

OpenStudy (anonymous):